OpenStudy (anonymous):

Solve for x: the quantity of x plus 16 all over 3 = 3x

OpenStudy (onepieceftw):

\[\frac{ x + 16 }{ 3 } = 3x\] Multiply both sides by 3 \[x + 16 = 9x\] Subtract x from both sides \[16 = 8x\] Divide both sides by 8 \[2 = x\]